OpenStudy (anonymous):

∠ RQS and ∠ TQS are a linear pair where m∠ RQS = 2x + 4 and m∠ TQS = 6x + 20 1. Solve for x. 2. Find m∠ RQS and m∠ TQS 3. Show how you can check your answer I want help to understand how to solve the problem because I have others to solve also.

OpenStudy (campbell_st):

linear pair means the angles add to 180 degrees so 2x + 4 + 6x + 20 = 180 collect like terms then solve for x

OpenStudy (anonymous):

2x + 4 + 6x + 20 = 180 -4 -4 = 2x +6x + 16 = 180

OpenStudy (anonymous):

do I have to subtract 4 from 180?

OpenStudy (campbell_st):

make it easier if you collect the like terms 1st you get 8x + 24 = 180 now its easier to solve

OpenStudy (anonymous):

ohh okay so then you do 8x + 24 = 180 -24 -24 = 8x = 156

OpenStudy (campbell_st):

yep thats right

OpenStudy (anonymous):

Then you do 8x/8 = 156/8 = x = 19.5

OpenStudy (campbell_st):

great... so you know x now you can find the measure of both angles bu=y substituting

OpenStudy (anonymous):

Thank you so much! :)